OpenStudy (anonymous):

Solve the system. y = x + 2 y = -x + 0 as and ordered pair x,y

OpenStudy (anonymous):

you can replace the first ecuation on the second so x+2 =-x 2x+2 =0 2x=-2 x= -1 then it on the other ecuation and y = -(-1) +0 y=1

OpenStudy (anonymous):

You could also add the 2 equations... y = x + 2 y = -x 2y = 2 y = 1 x = -1